One of the angles was formed on a straight line measured 93 degrees.What was the measure of its supplment?
I think the answer is 90 degrees.
But if anyone can help.Thank You!

Answer by TakeATuition.com(57)

SOLUTION -
WHEN SUM OF TWO ANGLES IS 180 DEGREE THEN THE ANGLES ARE CALLED AS SUPPLEMENTARY ANGLES.

ONE ANGLE IS 93
LET ANOTHER ANGLE BE X
SUM OF BOTH THE ANGLE IS 180
X+93=180
X+93-93=180-93
X= 87
FOR MORE HELP VISIT:www.TakeATuition.com